# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 3 LLM Diagnosis and Enhancement via Interpreting Model Components
This section discusses the XAI methods that interpret the internal components of large language models. Additionally, it delves into the insights these methodologies offer, which can be instrumental in refining and enhancing the design of language models. LLMs adopt transformers as the basic architecture, which typically comprises two types of major components: self-attention layers and feed-forward layers. In the following, we review the research that focuses on interpreting each of these components respectively.
3.1 Understanding the Self-Attention Module
A multi-head self-attention module contains multiple self-attention heads capturing different kinds of word-word relations, which are modeled with weights $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D_{1}\times D_{2}}$. Specifically, the relation of words $w_{i}$ and $w_{j}$ is computed as $\mathbf{A}_{i,j}\propto(\mathbf{x}_{i}\mathbf{W}_{q})\cdot(\mathbf{x}_{j}\mathbf{W}_{k})^{\top}$, where $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times D_{1}}$ are contextual embeddings of the words. The most straightforward interpretation is analyzing the attention score matrix $\mathbf{A}$ given an input sequence to study the relations between words (Vig, 2019; Hoover et al., 2020). In practice, these intuitive explanations would be majorly used to present case studies via visualization. With this strategy, Wang et al. (2023b) conduct case studies on in-context sentiment analysis, where they find that the label words from the in-context examples serve as anchors of the final prediction. In specific, these anchors aggregate information from the examples to generate informative representations from the lower layers, while the deeper layers utilize these representations to make final predictions. This insight motivates them to reweight the attention scores of these anchors to achieve better inference accuracy. Some researchers (Dar et al., 2023; Wu et al., 2023) extend this framework to globally analyze the attention weights $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ by feeding the static word embeddings of words from an interested vocabulary, instead of their contextual embeddings. For example, with this approach, Wu et al. (2023) find that instruction tuning empowers LLMs to follow human intentions by encouraging them to encode more word-word relations related to instruction words. On the other hand, some mathematical models are proposed to theoretically explain the self-attention mechanism, such as Sparse Distributed Memory (Bricken & Pehlevan, 2021) and Transformer Circuits (Elhage et al., 2021). In particular, the Transformer Circuits provide mechanistic interpretability to transformer-based models, which breaks models down into human-understandable pieces. Although these theoretical analyses on self-attention solids a foundation for future research, their direct application is largely underexplored.
3.2 Understanding the Feed-Forward Module
A feed-forward network is formalized as $\mathbf{x}^{\prime}=\sigma(\mathbf{x}\mathbf{W}_{u})\mathbf{W}_{v}^{\top}$, where $\mathbf{x}\in\mathbb{R}^{1\times D_{1}}$ is the intermediate contextual representation of an input word, $\sigma$ is a non-linear operation, and $\mathbf{W}_{u},\mathbf{W}_{v}\in\mathbb{R}^{D_{1}\times D_{3}}$ are model parameters. Feed-forward networks can be understood as key-value memories (Sukhbaatar et al., 2015; Geva et al., 2021), where each key or value is defined as $\mathbf{W}_{u}[d]\in\mathbb{R}^{D_{1}}$ and $\mathbf{W}_{v}[d]\in\mathbb{R}^{D_{1}}$, respectively. That is, each feed-forward network obtains $D_{3}$ key-value pairs, called memories. One simple way to interpret the semantic meaning of memory is collecting the words that could maximally activate the key or value vector of that piece of the memory (Geva et al., 2021; Dar et al., 2023), which has demonstrated strong interpretability of the extracted word lists. However, it is critical to be aware that the key or value vectors are polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023), indicating that this simple approach might not provide concise explanations for each key-value pair. It has been shown that the word list of each key-value pair has 3.6 human interpretable patterns on average (Geva et al., 2021). To alleviate the limited interpretability caused by the nature of polysemantic, Wu et al. (2023) propose to interpret the principal components of these key or value vectors, leading to a more concise explanation for each word list, such as “medical abbreviations” and “programming tasks and actions”. Other work examines individual memories by measuring the changes of predictions after perturbing their corresponding activations, where it reveals that some memories encode specific knowledge (Dai et al., 2022) and some others capture general concepts (Wang et al., 2022b). By leveraging the explanations of key-value memories, we could locate and update the memories associated with a specific piece of knowledge to perform model editing (Dai et al., 2022; Meng et al., 2022a; b; Hase et al., 2024), i.e., modifying outdated or incorrect knowledge. Another usage of these weight explanations is model pruning, where LLMs can be condensed up to 66.6% of their initial parameters by exclusively maintaining redundant neurons, resulting in an approximate 1.4x increase in inference speed (Dalvi et al., 2020).
Besides interpreting and analyzing model weights, some work studies model activations (e.g., $\mathbf{x}\mathbf{W}_{u}$ or $\mathbf{x}^{\prime}$) to interpret their functionality. The probing technique is the most popular way for this purpose (Belinkov et al., 2018; Tenney et al., 2018; Jawahar et al., 2019; Rogers et al., 2021), identifying whether a specific concept is encoded within the representations. The basic idea is developing an auxiliary classifier $g$ to map from the representations $\mathbf{x}^{\prime}$ to the interested concept space $\mathcal{C}$, such as syntax and part-of-speech knowledge, and the performance of $g$ interprets how much information encoded in $\mathbf{x}^{\prime}$ is related to the concepts in $\mathcal{C}$. This technique motivates developing better parameter-efficient (Chen et al., 2022), domain-specific (Das et al., 2023), and robust (Bai et al., 2021; Wang et al., 2023a) LLMs. Recent studies (Chen et al., 2023a; Ahdritz et al., 2024) also apply the probing method to detect the knowledge boundary of a LLM so that the hallucinated responses could be reduced. Some researchers (Bricken et al., 2023; Cunningham et al., 2023) point out another direction to interpret the model hidden activations, called dictionary learning, which is motivated by the assumption of superposition (Elhage et al., 2022; Sharkey et al., 2022). The superposition assumes that LLMs would learn an over-complete set of non-orthogonal features, thereby transcending the limitations imposed by the dimensionality of the representation space. Therefore, the researchers aim to reconstruct and interpret these features to understand the internals of the model. Practically, they develop a sparse auto-encoder $g$ to reconstruct the representations $\{\mathbf{x}_{n}\}$, which shows that humans could well interpret the learned sparse features of $g$ according to their most activation words. Their research shows that this method could be used for more controllable generation. Specifically, if forcing a sparse feature to be activated, then the language model $f$ would change its response to perform the particular behavior of that sparse feature. For example, given “1,2,3,4,5,6,7,8,9,10” as input, the model originally generates numbers as output. However, when they are forced to magnify the activations of a sparse feature called “DNA”, the model changes its output to “AGACCAGAGAGAAC”. In general, while the explanation techniques for feed-forward networks primarily offer insights for model development, they have also demonstrated promising applications in areas such as model editing and controllable generation.
3.3 Challenges
Interpreting the functionality of internal modules is still in its infancy, and we identify two challenges to be tackled in this direction.
3.3.1 Complexity of Individual Models and Their Interactions
The transformer-based language model contains two types of modules that collaborate based on the residual mechanism (He et al., 2016), which enables later modules to utilize, enhance, and/or discard outputs from preceding modules. Formally, the output of the $l$-th module is denoted as $x^{l}=f^{l}(x^{l-1})+x^{l-1}$, where $f^{l}$ could be a self-attention module or a feed-forward network. Research in this area aims to interpret how different modules $f^{i}$ and $f^{j}$ work together for $i\neq j$. Pilot studies (Elhage et al., 2021; Olsson et al., 2022) find that stacked self-attention modules could form Induction Heads, which demonstrate a strong correlation with the in-context learning capability. Specifically, the induction head encourages the model to predict the word “B” followed by a sequence “AB…A”. Their study finds a specific phase during pre-training LLMs where both induction heads and the in-context learning capability emerge from the model. Following this track, researchers observe diverse functional heads within LLMs for different tasks, such as “Name Mover Head” and “Duplicate Token Head” for the object identification task (Wang et al., 2022a), “Single Letter Head” and “Correct Letter Head” for the multiple-choice question answering task (Lieberum et al., 2023), and “Capitalize Head” as well as “Antonym Head” for the general purpose tasks (Todd et al., 2023). Although these studies have indeed deepened our understanding of cross-module effects, their analyses are grounded on specific tasks or scenarios, leaving it uncertain whether these findings possess broad generalizability.
3.3.2 Nature of Polysemantic and Superposition Assumption
Interpreting the functionality of a single neuron (one row/column vector of a weight matrix) fails in analyzing large language models since a single neuron could be activated by multiple and diverse meanings, called polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023). This nature leads to poor interpretability: explaining a single neuron usually does not reflect a concise human concept. Some researchers (Elhage et al., 2022; Sharkey et al., 2022) assume that this phenomenon is caused by the superposition of an over-complete set of features learned by the models. Based on this assumption, we may reach another level of explanation by decomposing the model weights to reconstruct a large number of features. However, the two critical problems of this approach are still unclear: (1) How do we ensure our reconstructed features faithfully represent the latent features learned by the models? (2) How do we interpret our reconstructed features with human language?
4 LLM Debugging with Sample-based Explanation

### v2

@@ 3. LLM Diagnosis and Enhancement via Interpreting Model Internals
This section discusses XAI methods that interpret LLM internals. We delve into the insights these methodologies offer, which can be instrumental in refining and enhancing the design of language models. Specifically, we review research that focuses on interpreting model components and latent representations, respectively, as shown in Figure 3.
3.1. Explanation of Model Components
LLMs adopt transformers as their basic architectures, which typically comprise two types of major components: Self-attention Mechanism (ATT) and Feed-forward Network (FFN). Researchers have focused on analyzing individual components as well as their interactions to gain insights into how LLMs make their predictions. We start by reviewing the connection between these components, focusing on how information is transformed through the residual stream. We then discuss methods to analyze the contribution of single or multiple components, respectively. We conclude by summarizing how these methods can be applied to improve model performance.
3.1.1. Linear Relation between Components
We consider a decoder-only LLM $f$ with $L$ internal layers, each of which is either an ATT or FNN layer, denoted as $f^{l}$. There is an input embedding matrix $\mathbf{W}_{\text{e}}$ that encodes each word of input text $x$ into a $D$-dimensional vector, i.e., $\mathbf{x}^{0}=\mathbf{W}_{\text{e}}[x]\in\mathbb{R}^{|x|\times D}$. The internal representation at each layer is computed by $\mathbf{x}^{l}=f^{l}(\mathbf{x}^{l-1})+\mathbf{x}^{l-1}$. Finally, there is an output embedding matrix $\mathbf{W}_{\text{u}}$ that decodes the hidden representations for next word prediction, i.e., $f(x)=\mathbf{x}^{L}\cdot\mathbf{W}_{\text{u}}$. By decomposing the final prediction (Elhage et al., 2021), we have
(2) $f(x)~{}=~{}\underbrace{\mathbf{x}^{0}\,\mathbf{W}_{\text{u}}}_{\text{Residual Stream}}~{}+~{}\sum_{l=1}^{L}\underbrace{f^{l}\Bigl{(}\mathbf{x}^{l-1}\Bigr{)}\,\mathbf{W}_{\text{u}}}_{\text{Layer Update}}.$
Equation (2) demonstrates two critical insights in understanding LLM mechanism: (1) Model prediction is a sum of layer outputs (Mickus et al., 2022). That is to say, the final prediction is an additive combination of the initial input embedding and the contribution from each layer. This linear transformation allows us to directly trace how each component shapes the overall prediction. Please note that the residual mechanism does not necessarily lead to a linear structure of model hidden spaces. For example, ResNet (He et al., 2016) does not show such a linear structure because there are non-linear operations applied to its residual stream. (2) The residual stream serves as a communication channel across layers (Elhage et al., 2021). In particular, each layer first extracts information from the residual stream, and then injects its processed output back into it. This shared residual stream enables multiple layers to work collaboratively, indirectly contributing to specific tasks by continuously updating the shared residual stream.
3.1.2. Importance Analysis on Single Component
Inspired by the first insight of Equation (2), early works (Geva et al., 2022; Ferrando et al., 2023; Wang et al., [n. d.]) quantify the contextual impact of a single layer $f^{l}$ in predicting a certain word $w$ under the context $x$ by computing $f^{l}(\mathbf{x}^{l-1})\,\mathbf{W}_{\text{u}}[:,w]$, where a greater value indicates a stronger impact. Following this trend, some researchers (Elhage et al., 2021; Meng et al., 2022a) propose to more precisely quantify the contextual impact by conducting interventions on layer outputs, i.e., $\text{diff}(\mathbf{h}\,\mathbf{W}_{\text{u}}[:,w],\widetilde{\mathbf{h}}\,\mathbf{W}_{\text{u}}[:,w])$, where $\text{diff}(\cdot,\cdot)$ measures the difference of two inputs, $\mathbf{h}=f^{l}(\mathbf{x}^{l-1})$, and $\widetilde{\mathbf{h}}$ is a intervention of $\mathbf{h}$. In practice, the choices of $\widetilde{\mathbf{h}}$ can vary, such as a constant vector (Wang et al., 2022a; Mohebbi et al., 2023), a noise vector (Meng et al., 2022a), and a hidden representation of other context (Hanna et al., 2023). On the other hand, researchers (Geva et al., 2021; Dar et al., 2023; Wu et al., 2023b) extend this framework to analyze the global impact of a single layer $f^{l}$ in predicting word $w$ by computing $\mathbf{W}^{l}_{\text{o}}\cdot\mathbf{W}_{u}[:,w]$, where $\mathbf{W}^{l}_{\text{o}}$ is the output weight matrix of layer $f^{l}$ that maps the processed results back to the residual stream.
3.1.3. Circuit Analysis on Multiple Components
Motivated by many emerging abilities of LLMs, it is interesting to explore how different layers work collaboratively to perform such complex tasks, called circuit analysis (Cammarata et al., 2020). Specifically, the goal of circuit analysis is to identify a subset of layers, whose combination performs a particular function, such as “indirect objects identification” (Wang et al., 2022a) and “number comparison” (Hanna et al., 2023). Recall that the residual stream enables the communication across layers, researchers (Geiger et al., 2021) formalize the forward pass of an LLM as a directed acyclic graph, where each node is one particular layer, and an edge is a representation flow from one layer to another via the residual stream. Under this setting, a circuit is a sub-graph with a distinct function. Identifying a specific circuit consists of three steps (Conmy et al., 2023): (1) selecting an interested task and constructing a dataset, (2) building the model’s graph, and (3) measuring the importance of sub-graphs with interventions (Sec. 3.1.2). This circuit identification process is time-consuming as it needs a large number of forward passes for interventions. To overcome this challenge, researchers (Sarti et al., 2023; Hanna et al., [n. d.]; Marks et al., 2024) propose a gradient-based method to approximate the intervention results with one more backward pass for each sample.
3.1.4. Usability of Component Analysis
The success of component analysis enables researchers to isolate the entire LLM into different parts for specific tasks. Therefore, the direct application of this technique is either modifying (enhancing or weakening) or removing certain parts that are essential or redundant to downstream tasks, resulting in two common applications: Knowledge Editing and Model Pruning.
  • Knowledge Editing. Component analysis can be used to locate the exact position storing a certain piece of knowledge (Dai et al., 2022; Wang et al., 2022b), and then update it accordingly to control the minimum effect on other knowledge. Specifically, we aim to update an outdated knowledge $t=(s,r,o)$ to a new one $t^{\prime}=(s,r,o^{\prime})$. Given the prefix $[s,r]$ of $t$, ROME (Meng et al., 2022a) first identifies a specific layer $f^{l*}\in f$ whose activation maximally contributes to predicting the original object $o$. This is done by maximizing the difference $f(o|[s,r])-f(o|[s,r],do(\mathbf{h}^{l}=\mathbf{h}^{l}+\epsilon))$, where $\mathbf{h}^{l}$ denotes the output activation for input $[s,r]$ at layer $f^{l}$, and $\epsilon$ denotes a random noise to simulate a perturbation. ROME then modifies only the parameters of this identified layer $f^{l*}$ to update the prediction to the new object $o^{\prime}$, without affecting unrelated knowledge stored elsewhere. The key limitation of ROME is that it can update only one piece of knowledge at a time, leading to an unacceptable time complexity to update a batch of knowledge. To overcome this challenge, follow-up works (Meng et al., 2022b; Hase et al., 2024; Li et al., 2024a; Gupta et al., 2024) scale up this process by skipping the knowledge locating step in ROME, and instead, they focus on learning the representation of updated knowledge at each layer. This line of work has been further applied to edit user personal information for privacy (Wu et al., 2023a), and edit out harmful concepts for safety (Wang et al., 2024c).
  • Model Pruning. Component analysis allows researchers to identify the specific roles of each model component for a given task, and thus, redundant or irrelevant components can be pruned to speed up the inference. Dalvi et al. (2020) directly measure the functionality of $f^{l}\in f$ by monitoring the expected difference $\mathbb{E}_{x,y\in\mathcal{D}}[p(y|x;f)-p(y|x,do(f^{l}(l)=\mathbf{0});f)]$, where $\mathcal{D}$ is a collected dataset for the downstream task. Recent studies (Men et al., 2024; Gromov et al., 2025) directly measure the redundancy by quantifying the similarities of hidden representations from different layers. Empirical studies (Dalvi et al., 2020; Fan et al., 2024) found that even over half of the model parameters are redundant for a certain task, and dropping them off can significantly reduce the inference time without sacrificing performance.
3.2. Explanation of Latent Representations
Many researchers have studied latent representations of LLMs (i.e., $\mathbf{x}^{l}$) to interpret model behaviors. It is critical to recognize that the representations are not naturally interpretable because of their polysemantic nature (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023), indicating that each dimension of the latent space represents multiple semantic meanings, called concepts. Thus, we cannot clearly understand a latent representation by reviewing the values in each dimension. There are various lines of work to tackle this challenge, and we categorize them according to whether they require additional training.
3.2.1. Non-Training based Methods
Both the input and output layer of transformer-based LLM $f$ are directly interpretable as they operate on human-readable text. Hidden representations across different layers share the same residual stream (see Section 3.1.1), allowing information to flow seamlessly between layers. Consequently, non-training based interpretation methods leverage this insight by transferring the hidden representation $\mathbf{x}^{l}$ to either the first or the last layer to facilitate interpretation.
  • Projecting $\mathbf{x}^{l}$ to Output Embeddings. Given a hidden representation $\mathbf{x}^{l}$ at layer $l$, Logit Lens (nostalgebraist, 2020) understands its semantics by projecting it to the output word embedding matrix $\mathbf{W}_{\text{u}}$. Specifically, we interpret $\mathbf{x}^{l}$ by collecting the $K$ words whose output embeddings could maximally activate $\mathbf{x}^{l}$, i.e.,
    (3) $\mathcal{I}=\underset{\mathcal{V}^{\prime}\subset\mathcal{V},|\mathcal{V}^{\prime}|=K}{\arg\max}\sum_{w\in\mathcal{V}^{\prime}}\mathbf{x}^{l}\cdot\mathbf{W}_{\text{u}}[:,w],$
    where $\mathcal{V}$ is a pre-defined vocabulary set. The Logit Lens can be viewed as an instance of skipping all later layers after $\mathbf{x}^{l}$ (Jastrzebski et al., 2018). However, researchers (Belrose et al., 2023; Din et al., 2024) find that there is a distribution shift between the intermediate layer $l$ and the last layer $L$, and thus, they propose to train a translator to learn the distribution shift to improve its interpretability for representations from shallow layers.
  • Feeding $\mathbf{x}^{l}$ as Input Embeddings. In contrast with moving hidden representation $\mathbf{x}^{l}$ to the output layers, Self-IE (Chen et al., 2024b) proposes feeding $\mathbf{x}^{l}$ as an input word embedding in another round of model inference, where we prompt LLM to explain the meaning of the fed representation. For example, we may instruct LLM to interpret $\mathbf{x}^{l}$ with the prompt “Please summarize [X] in one sentence.”, where “[X]” is a placeholder, whose embedding will be replaced with our interested hidden representation $\mathbf{x}^{l}$. Since LLMs are trained to follow human instructions, they are expected to generate a sentence to describe the information encoded in $\mathbf{x}^{l}$.
3.2.2. Supervised Training based Methods
The supervised training-based interpretation methods break the polysemantic nature of LLM latent space by leveraging some annotated datasets. They focus on exploring two main research questions: (1) whether the LLM latent space encodes a specific concept; and if so, (2) whether this concept impacts LLM’s predictions and how significant.
  • Identifying Existence of Certain Concepts. Probing (Gupta et al., 2015) is the most traditional technique in analyzing the knowledge encoded within hidden representations of LLMs. In specific, the Probing technique requires an annotated dataset to train a probe $p$ with parameter $\theta$ to predict the occurrence of interested knowledge, i.e., $p:\mathbb{X}\rightarrow[0,1]$, where $\mathbb{X}$ is the latent space of LLMs at a certain layer. Here, the dataset consists of positive samples (with label “1”) that clearly demonstrate our interest concept and negative samples that clearly exclude such concept. For example, early works (Belinkov et al., 2017) use this technique to identify some morphology concepts (e.g., part-of-speech tags and morph tags) in the LLM’s latent space by constructing a dataset where each example has human annotations on these tags. If the probe $p$ achieves a significantly high accuracy on a certain dataset, then we could conclude that the latent representation encodes that concept. This technique has been used to study diverse concepts, ranging from low-level linguistic patterns (Belinkov et al., 2017; Hewitt and Manning, 2019; Chen et al., 2021; Liu et al., 2019; Hewitt and Manning, 2019) to high-level semantic meanings (Burns et al., 2022; Zou et al., 2023a; MacDiarmid et al., 2024; Zheng et al., 2024). The improvements in the probing technique (Hewitt and Liang, 2019; Pimentel et al., 2020; Voita and Titov, 2020) mainly focus on designing baselines to confirm that the high accuracy achieved by probe $p$ is led by the knowledge provided by the latent space $\mathbf{x}$.
  • Measuring Importance of Certain Concepts. To quantify how much a specific concept impacts the model’s predictions, researchers often use TCAV (Kim et al., 2018). We first train a linear probe $p$ with parameter $\theta\in\mathbb{R}^{D}$ to distinguish latent representations containing concept $c$ or not, where $\theta$ is named as the concept activation vector (CAV). TCAV measures how sensitive the model’s output is to perturbations along this concept direction (e.g., via directional derivatives of the model’s logits). Intuitively, if the model’s predictions change substantially when the representation is shifted along the learned concept direction, we can conclude that the concept has a high importance for the model’s decision. Mathematically, the perturbation is realized by computing the gradient of $f(x)$ with respect to $\mathbf{x}^{l}$ and projecting it onto $\theta$ as $\Delta=\nabla_{\mathbf{x}^{l}}f(x)\cdot\theta$, so that a larger $|\Delta|$ indicates that small shifts in the latent space along $\theta$ lead to substantial changes in $f(x)$, underscoring the concept’s influence.
3.2.3. Unsupervised Training based Methods
Breaking the polysemantic constraint for interpretation can be formalized as an unsupervised learning problem on the basis vectors of the LLM latent space (Fel et al., 2023), where each basis vector refers to a clear and concise semantic concept, indicating a monosemantic nature. Mathematically, given a total of $N$ hidden representations $\mathbf{X}\in\mathbb{R}^{N\times D}$, we aim to learn two matrices $\mathbf{A}\in\mathbb{R}^{N\times C}$ and $\mathbf{C}\in\mathbb{R}^{C\times D}$ by minimizing the loss
(4) $\underset{\mathbf{A},\mathbf{C}}{\arg\min}\,\,\|\mathbf{X}-\mathbf{A}\mathbf{C}\|^{2},$
where $\mathbf{C}$ serve as the learned concepts, and $\mathbf{A}$ is the coefficient of the instances on each concept. There are some additional conditions to ensure that the learned concept vectors satisfy our expectations on monosemantic.
  • Orthogonal Assumption on Concepts. Early works (Millidge and Black, 2022; Wu et al., 2023b) applies the Singular Vector Decomposition to learn a set of concept vectors $\mathbf{C}$ that are perfectly orthogonal, i.e.,
    (5) $\mathbf{C}\mathbf{C}^{\top}=\mathbf{I},$
    where $\mathbf{I}$ is an identical matrix. However, satisfying this hard request is computing costly in practice, leading to limited scalability in interpreting hidden representations from giant LLMs with both large $N$ and $D$.
  • Sparse Assumption on Concepts. Sparse Autoencoder (Olshausen and Field, 1997) (SAE) is another alternative that relaxes the orthogonal condition to be a smoother one in which the sample coefficients matrix is asked to be sparse, i.e.,
    (6) $\mathbf{A}=\text{Top-K}(\mathbf{X}\cdot\mathbf{C}^{\top}),$
    where Top-K activation enforces all values to be zero unless the $K$ largest ones from the inputs. Specifically, researchers (Cunningham et al., 2023; Bricken et al., 2023) train a SAE $g(\mathbf{x})=\text{Top-K}(\mathbf{x}\cdot\mathbf{W}^{\top})\cdot\mathbf{W}$ to reconstruct hidden representations $\mathbf{x}$ by minimizing $\|\mathbf{x}-g(\mathbf{x})\|^{2}$, where $\mathbf{W}\in\mathbf{R}^{C\times D}$ and Top-K activation enforces all values to be zero unless the $K$ largest ones from the input. This approach has shown effectiveness in decomposing the latent representations from giant LLMs with a hundred billion parameters, such as GPT-4 (Gao et al., 2024) and Claude 3 (Templeton et al., 2024).
Once these concept vectors are obtained, their representing meanings can be further described with natural language by selecting $M$ input texts that could maximally activate the related concept vectors (Bricken et al., 2023), i.e.,
(7) $\mathcal{I}_{c}=\arg\max_{\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|=M}\sum_{x\in\mathcal{X}^{\prime}}f^{l}(x)\mathbf{C}[c].$
However, the latest research (Wu et al., 2025c; Gur-Arieh et al., 2025) has shown that such an input-based explanation cannot well interpret their impacts on controlling LLM responses, and thus, they propose to interpret these learned concept vectors by building the connection to the output texts. We refer audiences to check (Shu et al., 2025) for a more detailed review of designing and interpreting SAEs for LLMs.
3.2.4. Usability of Interpreting Latent Representations
Latent space interpretations improve LLMs further at each stage of their entire life-cycles. In particular, we consider three stages of model development, including the “pre-processing” stage for data preparation, the “training” stage that trains LLM on prepared data, and the “inference” stage that uses a trained model to operate user requests.
  • Data Pre-processing. Existing studies have empirically found that the same model but trained on different datasets with different diversity (Ge et al., 2024), complexity (Xu et al., 2023c), and quality (Li et al., 2024c) levels may lead different downstream task performance. The latent space interpretation techniques enable researchers to understand the geometric structure of a dataset and, further, to select a subset that preserves a certain geometric structure. For example, to obtain a subset of the dataset preserving as diverse topics as possible, SAE-GreedSelect (Yang et al., 2025b) greedily select examples to maximize the total number of activated SAE features. Results show that a selected subset with only 5% of total samples can train a model with comparable performance to a full-trained one, significantly reducing the training cost.
  • Training-Time Regularization. Either the trained parameter $\theta$ of probe $p$ or each column vector $\mathbf{C}[c]$ learned by SAE or SVD indicates a feature in the LLM latent space. Therefore, these learned feature vectors can be used to perform feature engineering in the LLM latent space for model training. For example, Yin et al. (2024a) propose a feature-level constrained fine-tuning approach. This method effectively prevents the fine-tuned model from changing significantly from its pretrained counterpart without explicitly involving the pretrained model during fine-tuning, thereby substantially reducing computational costs. In contrast, Wu et al. (2025b) introduce a more semantically meaningful regularization approach for model training. They first determine whether each learned feature vector $\mathbf{C}[c]$ is causally relevant to the downstream task by analyzing its natural language explanation $\mathcal{I}_{c}$. Subsequently, they impose a penalty during fine-tuning to discourage reliance on features identified as task-irrelevant (i.e., shortcuts). This approach effectively enhances the generalizability of the fine-tuned model.
  • Inference-Time Control. As the learned feature vectors by probing or SAEs break the polysemantic nature and achieve monosemantic, it is possible to directly steer LLM latent space toward a specific direction without sacrificing other abilities. Formally, let $\mathbf{z}$ denote the learned feature vector to an interested concept, we can erase it from the hidden representation of input $x$ by $\mathbf{x}^{\prime}=\mathbf{x}-\text{ReLU}(\mathbf{x}\cdot\mathbf{z}^{\top})\cdot\mathbf{z}$, or constantly amplify it to a certain level $\alpha$ by $\mathbf{x}^{\prime}=\mathbf{x}+\alpha\cdot\mathbf{z}$. The combination of these two techniques on different categories of feature vectors leverages model steering for different purposes. For example, Wu et al. (2025c) observe that certain feature vectors semantically correlate with specific safety-related concepts, such as “First Aid,” based on summarizing their natural language explanations $\mathcal{I}_{c}$. By enforcing a constant activation of these safety-related feature vectors during model generation, they demonstrate that the model reliably switches from generating harmful responses to explicitly rejecting inappropriate or malicious user inputs. Generally, this technique has shown success in steering LLMs during inference to improve their helpfulness (Huang, 2024; Stolfo et al., 2024; Rimsky et al., 2024; He et al., 2025), safety (Arditi et al., 2025; Bhattacharjee et al., 2024; Zou et al., 2023a; Rimsky et al., 2024), faithfulness (Chen et al., 2023a; Ahdritz et al., 2024), fairness (Singh et al., 2024; Zou et al., 2023a), and currentness (Scialanga et al., 2025).
3.3. Challenges
3.3.1. Explanation Efficiency
The most significant challenge in analyzing the inner workings of LLMs is their demand for a large amount of computing resources. Circuit analysis (Section 3.1.3) is one of the examples that is bounded by this bottleneck. Specifically, performing circuit analysis on one sample still requires a large amount of forward inference of LLMs under different perturbed internals of that sample, leading to limited scalability to analyze over a large dataset. Therefore, existing works on circuit analysis focus on studying particular behaviors of LLMs, such as “Name Mover Head” and “Duplicate Token Head” for object identifications (Wang et al., 2022a), “Single Letter Head” and “Correct Letter Head” for multiple-choice question answering (Lieberum et al., 2023), and “Capitalize Head” as well as “Antonym Head” for writing (Todd et al., 2023). On the other hand, although SAEs (Section 3.2.3) have shown strong promise in understanding LLM latent representations, training such proxy models requires collecting LLM representations over a large corpus. For example, researchers (Gao et al., 2024) found that training an SAE with only 1 million features requires over 10 billion training tokens. Future works may explore computing-efficient methods to understand LLMs’ diverse behaviors comprehensively.
3.3.2. Explanation Faithfulness
Although many proposed methods have shown strong promise with certain usability cases (Section 3.1.4 and Section 3.2.4), some researchers still hold concerns about whether they faithfully interpret the inner workings of LLMs since they may not practically outperform some trivial baselines. For example, although many researchers (Wu et al., 2025c; Bricken et al., 2023) can apply SAEs (Section 3.2.3) to steer LLM responses for specific purposes, some research (Wu et al., 2025a; Durmus et al., 2024; Bricken et al., 2024) has observed that SAE-based methods cannot significantly outperform prompting baselines, posing a serious concern about the effectiveness of SAEs. Furthermore, other researchers (Heap et al., 2025; Paulo and Belrose, 2025) suspect that SAEs cannot faithfully explain LLMs’ internals, as they observe that SAEs can even learn features from randomized LLMs. These concerns suggest two critical questions for future works: (1) How do we ensure our learned features faithfully represent LLM internals? (2) In which cases are SAE-based methods more effective than prompting strategies?
4. LLM Debugging with Sample-based Explanation